on 1st January 2017, Mwanawasa Ltd, had in issue. 800.000 Ordinary shares with a par value of sh 10 each. on 31st october 2017 the company issue 100,000 Ordinary shares at full market Price of Sh 20 cach. on 41st December 2018, the director decided to declare a bonus issue of I for 4 on 31st December 2019, thedirectors decided to declare a a right issue for 1 for 5 at a Price of sh 20. on the day immediately before for right issue took effect, the shares of this company were trading qt sh 25 each. Mwanawasa Itd does not have any preference Share in issue. The Post tax earning for the year 2017, 2018 1 and 2019 were £ 350.000 and € 520.000 and £ 430.000 respectively.


Required


a) compute the Basic EPS for the years 2017, 2018, 2019.



b) comment on the financial Performance of the company in the year 2018 and 2019.

1
Expert's answer
2022-03-09T10:31:18-0500

a)

"EPS= \frac{End-of-Period Common Shares Outstanding}{Net Income − Preferred}"

2017= "\frac{£ 350,000}{100,000}= 3.5"

2018= "\frac{£520,000}{125,000}= 4.16"

2019= "\frac{£430,000}{300,000}= 1.433"

b) The performance of the organization was not good as the ratio went down.